The length of a rectangle is 4 centimeters longer than the width of the rectangle. The perimeter of the
rectangle is 88 centimeters. What is the area of the rectangle, in square centimeters?

Answer by fractalier(6550)

Perimeter P = 2L + 2W = 88
We also have L = W + 4
Now substitute this into the first equation and get
2(W + 4) + 2W = 88
2W + 8 + 2W = 88
4W + 8 = 88
4W = 80
W = 20 cm
L = 24 cm